# CLMFormer: Mitigating Data Redundancy to Revitalize Transformer-based Long-Term Time Series Forecasting System

### v4

@@ IV-E Ablation Study
The ablation study detailed in the Table.V examines the impact of various components of the proposed model on time series forecasting accuracy across three datasets: Etth1, Air Quality, and Traffic. This study specifically looks at the performance of the Informer and FedFormer model and their enhancements through progressive training (p.t.), our memory approach (mem), and a combination of both referred to as “ours.” From the results, we can draw several key observations. Both the progressive training strategy and the memory approach independently improve the Informer’s forecasting accuracy. For instance, in the Etth1 dataset, applying progressive training reduced the MSE from 1.110 to 1.021 for a 168-hour prediction horizon, and incorporating the memory approach further reduced it to 1.009. Similar improvements are seen in MAE metrics and across other datasets. The combination of progressive training and the memory approach under“Informer + ours” consistently yields the best results. For example, it achieves the lowest MSE of 0.495 and MAE of 0.493 for a 24-step forecast on the Etth1 dataset. This pattern holds across all datasets and prediction horizons, indicating a synergistic effect when both enhancements are applied together. The improvements are more pronounced for longer prediction horizons. For instance, in the Traffic dataset, the ”Informer + ours” configuration reduces the MSE to 0.405 and the MAE to 0.300 for a 168-step forecast, suggesting that the proposed components are particularly effective at addressing challenges in long-term sequence forecasting (LTSF). While the Informer model and its enhancements focus on incremental improvements, the FedFormer model, both alone and with added enhancements (progressive training, memory approach, and combined), shows varied performance across datasets. “FedFormer+ours” achieves better MSE and MAE scores in the shorter and longer prediction horizons for the Etth1 dataset, and always outperforms the base model, especially in the Air Quality and Traffic datasets for longer horizons. These detailed observations confirm the effectiveness of the combined approach in improving forecasting accuracy, especially in longer prediction horizons. This suggests that the integration of these components can effectively tackle the inherent challenges in LTSF problems, making it a valuable strategy for enhancing time series forecasting models.
@@
Table VI presents the findings from our second ablation study focused on evaluating the efficacy of different memory matrix initialization methods within our proposed model across three distinct datasets: Etth1, Etth2, and Weather. This study specifically compares the performance of our model using a seasonal memory matrix (denoted as ”ours”) against a baseline initialization approach employing an Identity matrix [45, 46]. The results delineated in the table underscore the superior performance of the Informer model augmented with our seasonal memory matrix across the majority of tasks and horizons, particularly highlighting its robustness in capturing long-term dependencies and complex temporal dynamics inherent in the datasets. For instance, in the Etth1 dataset, our approach consistently outperforms the Identity matrix initialization in both MSE and MAE metrics across all forecasting horizons, with significant improvements observed at longer horizons (336 and 720 predictions), which underscores the enhanced capacity of our model to retain and leverage long-term historical information for accurate future predictions. Moreover, the study also explores the efficacy of these memory matrix initialization methods within the FedFormer model, further substantiating the adaptability and effectiveness of our seasonal memory matrix in enhancing forecasting performance. Notably, the FedFormer model with our memory matrix initialization demonstrates remarkable improvements over the Identity matrix, especially in the Etth2 and Weather datasets, indicating the generalizability of our proposed method across different models and datasets. This ablation study clearly demonstrates that our seasonal memory matrix significantly outperforms the traditional Identity matrix in capturing long-term dependencies and intricate temporal patterns within time series data. By providing a more sophisticated mechanism for initializing the memory matrix, our approach enables the model to better understand and predict future values based on historical data, thereby setting a new benchmark for performance in time series forecasting tasks. This validation not only reinforces the value of our adaptations to the memory matrix but also highlights the potential for further innovations in enhancing the capabilities of time series forecasting models.
IV-F Case Study
@@
In this section, we embark on two case studies to conduct a qualitative evaluation of the predictive capabilities of our CLMFormer and FedFormer. Specifically, we present the input and ground truth instances with the predictions rendered by both FedFormer and CLMFormer over a forecast length of 720 intervals within the ETTh1 and Air Quality datasets, as depicted in Figure 4. Across both subplots, it is discernible that the predictions generated by our CLMFormer, represented in green, more closely adhere to the ground truth, delineated in blue, exhibiting similar trends in contrast to those of FedFormer. This convergence with the ground truth underscores the enhanced capability of CLMFormer to generate both extensive and high-fidelity predictions. Moreover, it becomes evident that the FedFormer’s predictions, illustrated in purple, tend to perpetuate the trends observed in the input data, adhering to a linear trajectory. In contrast, CLMFormer is adept at capturing non-linear relationships. We postulate that this is attributable to our proposed seasonal memory-driven forecasting modules, which empower the model to internalize seasonal patterns from the entire corpus of training data, thereby furnishing more accurate and elongated predictive insights.
V Conclusion
@@
In this paper, we propose a novel pipeline, named CLMFormer, to revitalize Transformer-based LTSF systems by mitigating data redundancy stemming from the rolling forecasting settings. We first discuss the influence of rolling forecasting settings on Transformer-based LTSF systems. Subsequently, we train a seasonal memory-driven Transformer-based model in a progressive fashion. We demonstrate that seasonal memory improves the model’s ability to comprehend both multivariate and univariate sequential patterns from highly similar training samples. In addition, the progressive training schedule can effectively address overfitting issues in LTSF. We validate our methods on six public real-life benchmarks with five different prediction lengths ranging from 24 to 720 time steps. The results show that our method works exceptionally well with longer prediction sequences and can be seamlessly integrated into varying Transformer-based LTSF models and achieve up to 30% improvement.

### v5

@@ 4.5. Ablation Study
The ablation study is detailed in the Table.5 examines the impact of various components of the proposed model on time series forecasting accuracy across three datasets: Etth1, Air Quality, and Traffic. This study specifically looks at the performance of the Informer and FedFormer model and their enhancements through progressive training (p.t.), our memory approach (mem), and a combination of both referred to as “ours." From the results, we can draw several key observations. Both the progressive training strategy and the memory approach independently improve the Informer’s forecasting accuracy. For instance, in the Etth1 dataset, applying progressive training reduced the MSE from 1.110 to 1.021 for a 168-hour prediction horizon, and incorporating the memory approach further reduced it to 1.009. Similar improvements are seen in MAE metrics and across other datasets. The combination of progressive training and the memory approach under“Informer + ours" consistently yields the best results. For example, it achieves the lowest MSE of 0.495 and MAE of 0.493 for a 24-step forecast on the Etth1 dataset. This pattern holds across all datasets and prediction horizons, indicating a synergistic effect when both enhancements are applied together. The improvements are more pronounced for longer prediction horizons. For instance, in the Traffic dataset, the ”Informer + ours" configuration reduces the MSE to 0.405 and the MAE to 0.300 for a 168-step forecast, suggesting that the proposed components are particularly effective at addressing challenges in long-term sequence forecasting (LTSF). While the Informer model and its enhancements focus on incremental improvements, the FedFormer model, both alone and with added enhancements (progressive training, memory approach, and combined), shows varied performance across datasets. “FedFormer+ours” achieves better MSE and MAE scores in the shorter and longer prediction horizons for the Etth1 dataset, and always outperforms the base model, especially in the Air Quality and Traffic datasets for longer horizons. These detailed observations confirm the effectiveness of the combined approach in improving forecasting accuracy, especially in longer prediction horizons. This suggests that the integration of these components can effectively tackle the inherent challenges in LTSF problems, making it a valuable strategy for enhancing time series forecasting models.
Figure 5. Samples of time series data from the ETTh1, Weather, and Air Quality datasets, along with predictions by FedFormer and our CLMFormer. We employ multivariate settings with prediction lengths of 720 and visualize the OT, temperature and PM10 value for the ETTh1, Weather and Air Quality datasets, respectively.
@@
Table.6 presents the findings from our second ablation study focused on evaluating the efficacy of different memory matrix initialization methods within our proposed model across three distinct datasets: Etth1, Etth2, and Weather. This study specifically compares the performance of our model using a seasonal memory matrix (denoted as "ours") against a baseline initialization approach employing an Identity matrix (chen2020raio, ; memoryma2021, ). The results delineated in the table underscore the superior performance of the Informer model augmented with our seasonal memory matrix across the majority of tasks and horizons, particularly highlighting its robustness in capturing long-term dependencies and complex temporal dynamics inherent in the datasets. For instance, in the Etth1 dataset, our approach consistently outperforms the Identity matrix initialization in both MSE and MAE metrics across all forecasting horizons, with significant improvements observed at longer horizons (336 and 720 predictions), which underscores the enhanced capacity of our model to retain and leverage long-term historical information for accurate future predictions. Moreover, the study also explores the efficacy of these memory matrix initialization methods within the FedFormer model, further substantiating the adaptability and effectiveness of our seasonal memory matrix in enhancing forecasting performance. Notably, the FedFormer model with our memory matrix initialization demonstrates remarkable improvements over the Identity matrix, especially in the Etth2 and Weather datasets, indicating the generalizability of our proposed method across different models and datasets. This ablation study clearly demonstrates that our seasonal memory matrix significantly outperforms the traditional Identity matrix in capturing long-term dependencies and intricate temporal patterns within time series data. By providing a more sophisticated mechanism for initializing the memory matrix, our approach enables the model to better understand and predict future values based on historical data, thereby setting a new benchmark for performance in time series forecasting tasks. This validation not only reinforces the value of our adaptations to the memory matrix but also highlights the potential for further innovations in enhancing the capabilities of time series forecasting models.
4.6. Case Study
@@
In this section, we embark on two case studies to conduct a qualitative evaluation of the predictive capabilities of our CLMFormer and FedFormer. Specifically, we present the input and ground truth instances with the predictions rendered by both FedFormer and CLMFormer over a forecast length of 720 intervals within the ETTh1, Weather and Air Quality datasets, as depicted in Figure.5. Across both subplots, it is discernible that the predictions generated by our CLMFormer, represented in green, more closely adhere to the ground truth, delineated in blue, exhibiting similar trends in contrast to those of FedFormer. This convergence with the ground truth underscores the enhanced capability of CLMFormer to generate both extensive and high-fidelity predictions. Moreover, it becomes evident that the FedFormer’s predictions, illustrated in purple, tend to perpetuate the trends observed in the input data, adhering to a linear trajectory. In contrast, CLMFormer is adept at capturing non-linear relationships. We postulate that this is attributable to our proposed seasonal memory-driven forecasting modules, which empower the model to internalize seasonal patterns from the entire corpus of training data, thereby furnishing more accurate and elongated predictive insights.
5. Limitations and Future Work
5.1. Limitations
One limitation of CLMFormer is its reliance on predefined curriculum learning schedules, where noise injection follows a fixed monotonic function. This approach may not always be optimal across different datasets, as the ideal noise level may vary based on the characteristics of each dataset. Additionally, while the memory-driven decoder enhances the model’s ability to capture seasonal dependencies, its fixed update mechanism may not be flexible enough for highly dynamic environments where trends shift rapidly. Lastly, CLMFormer, like most Transformer-based models, still requires a substantial amount of training data to generalize well. Although our curriculum learning strategy helps mitigate overfitting, performance on extremely small datasets remains a challenge.
5.2. Future Work
Several promising directions can further improve CLMFormer. First, adaptive curriculum learning could replace the fixed noise injection strategy by incorporating reinforcement learning or meta-learning to dynamically adjust the dropout rate based on dataset characteristics, ensuring optimal training progression. Second, the memory-driven mechanism could be extended to include personalized memory modules, allowing for adaptive seasonal pattern modeling tailored to specific time-series sequences, making it particularly beneficial for applications such as patient-specific healthcare forecasting or financial market predictions. Third, scaling CLMFormer to large-scale foundation models could enable pretraining on diverse multi-domain time-series datasets, improving generalization and enabling cross-domain transfer learning. Addressing these challenges will further enhance the flexibility, efficiency, and adaptability of CLMFormer in real-world forecasting applications.
6. Conclusion
@@
In this paper, we propose a novel pipeline, named CLMFormer, to revitalize Transformer-based LTSF systems by mitigating data redundancy stemming from the rolling forecasting settings. We first discuss the influence of rolling forecasting settings on Transformer-based LTSF systems. Subsequently, we train a seasonal memory-driven Transformer-based model in a progressive fashion. We demonstrate that seasonal memory improves the model’s ability to comprehend both multivariate and univariate sequential patterns from highly similar training samples. In addition, the progressive training schedule can effectively address overfitting issues in LTSF. We validate our methods on six public real-life benchmarks with five different prediction lengths ranging from 24 to 720-time steps. The results show that our method works exceptionally well with longer prediction sequences and can be seamlessly integrated into varying Transformer-based LTSF models and achieve up to 30% improvement.